Use the Method of compensation to show how you can make the subtraction of 563 -79 easier


1
Expert's answer
2021-05-31T16:31:54-0400

Subtraction by compensation means taking away more than you need and then adding some back. It can be quite an effective way to subtract.

Now:


"\Rightarrow" "563-79"

"\Rightarrow" "563-100+21" (break -79=-100+21,as it is easy to subtract 100. )

It can easily be:

"\Rightarrow" "463+21"

"\Rightarrow463+20+1=484"


484 is the real answer ,which we have determine so easily by the help of compensation method.